OpenStudy (anonymous):

if there is 1176 handshakes ..then how many members are present?

OpenStudy (saifoo.khan):

1176/2

OpenStudy (anonymous):

That is only true if every one only handshakes one person...

OpenStudy (anonymous):

Yeah, I'd say the question is incomplete

OpenStudy (saifoo.khan):

Yeah!

OpenStudy (anonymous):

nope, lets use the formula for a Sum of numbers Sum= n(n+1)/2 = 1076 and solve for n \[n ^{2}+n -2352 = 0\] the two possible values of x are (using the quadratic equation formula= x = 48 & x= - 49 And since there cannot be negative amount of people. There are 48 people

OpenStudy (anonymous):

basically here, let us put all the people in a line and the last one will shake hands with everyone and then will leave the room... ok? the following person will shake all but the last one (n-1) and so on until there is only one person, which will not shake anyones hand, as he has already done it with every one. This a basic sum \[{\sum_{0}^{x}} = {n (n+1) \over 2}\]